Gunderson Consulting provides a wide range of consulting services for major league Baseball teams.  They are considering different methods of charging costs to individual jobs in a job order costing system.  The majority of the billable services are between the world series and beginning of spring training.  The support workers complete research and analysis evenly throughout the year except they all get the month of December off.   For the last year they incurred total direct costs of $400,000 and 10,000 direct billable hours.  Before the year began they had not expected the pandemic, so they had budgeted $1,000,000 direct costs and budgeted 20,000 billable hours.  Overhead had been budgeted $500,000 and incurred $350,000.  They allocate the overhead based on billable hours.

  1. Calculate the cost per billable hour for direct costs and indirect costs under:
    1. Actual costing
    2. Normal costing
    3. Variation of normal costing
    4. Standard costing
  2. Determine the amount of costs charged to a job that incurred 50 hours, but had budgeted 60 hours under:
    1. Actual costing
    2. Normal costing
    3. Variation of normal costing
    4. Standard costing
  3. Which of the different costing methods would you recommend?  State your answer in paragraph form (not bullets).  Explain your answer.
Answer #1

Given information

Total acutal costs

total direct costs of $400,000

direct billable hours 10000

Actual Overhead   $350,000

Total budgeted costs

budgeted direct costs $1,000,000

budgeted billable hours 20000

Overhead budgeted $500,000

note They allocate the overhead based on billable hours.

particulars actual normal standard
direct labour actual price * Actual qty actual price * Actual qty Standard price * standard qty
variable overhead actual standard rate * actual qty standard rate * standard qty

SP (SR): standard price (rate for overhead) is the budgeted price (rate).

SQ: standard quantity is the budgeted quantity

1. Calculate the cost per billable hour for direct costs and indirect costs under:

direct cost

1 standard price : 1000000/20000 = 50 per hr

2 standard qty : 20000

3 actual price : 400000/10000 = 40 per hr

4 actual qty : 10000

overhead (indirect) cost

1 standard price : 500000/20000 = 25 per hr

2 standard qty : 20000

3 actual price : 350000/10000 = 35 per hr

4 actual qty : 10000

particulars actual normal standard
direct labour actual price * Actual qty actual price * Actual qty Standard price * standard qty
40*10000 = 400000 40* 10000 = 400000 50 * 20000 = 1000000
variable overhead actual standard rate * actual qty standard rate * standard qty
350000 25 * 10000 = 250000 25* 20000 = 500000
  1. Actual costing : 400000+350000 = 750000
  2. Normal costing : 400000+250000 = 650000
  3. Variation of normal costing : 750000-650000 = 100000
  4. Standard costing : 1000000+500000 = 1500000

2. Determine the amount of costs charged to a job that incurred 50 hours, but had budgeted 60 hours under:

direct cost

1 standard price : 1000000/20000 = 50 per hr

2 standard qty : 60

3 actual price : 400000/10000 = 40 per hr

4 actual qty : 50

overhead (indirect) cost

1 standard price : 500000/20000 = 25 per hr

2 standard qty : 60

3 actual price : 350000/10000 = 35 per hr

4 actual qty : 50

particulars actual normal standard
direct labour actual price * Actual qty actual price * Actual qty Standard price * standard qty
40*50 = 2000 40* 50 = 2000 50 * 60 = 3000
variable overhead actual standard rate * actual qty standard rate * standard qty
35 * 50 = 1750 25 * 50 = 1250 25* 60 = 1500
  1. Actual costing : 2000+1750 = 3750
  2. Normal costing : 2000+1250 = 3250
  3. Variation of normal costing : 3750-3250 = 500
  4. Standard costing : 3000+1500 = 4500

3.Which of the different costing methods would you recommend?  State your answer in paragraph form (not bullets).  Explain your answer.

normal costing is recomended because the cost is low and as well the profits will raise. so normal costing is considered better

kindly upvote for me, it will help me alot, thank you!!